Calculate the total change in aggregate spending if investment increases by $200 billion and the marginal propensity to consume is 0.8.
Answer #1

Answer

Multiplier =1/(1-MPC)

=1/(1-0.8)

=5

total change in aggregate spending =change in investment *multiplier

=200*5

=$1000

the total change in aggregate spending is $1000 billion